OpenStudy (maheshmeghwal9):

Please convert this into trigonometrical form: - \[\sin 32^o+i \cos 32^o.\]

OpenStudy (shubhamsrg):

assuming you meant exponential form : we have this : -i^2 sin32 + icos32 take i common : i(cos32 - isin32) =ie^(-i32)

OpenStudy (experimentx):

\[ \huge a + ib = e^{\ln |\sqrt{a^2 + b^2 }| + \tan^{-1}\left ( b \over a\right )}\]

OpenStudy (experimentx):

and covert that degree into radians ..

OpenStudy (maheshmeghwal9):

But In my book answer is given as wait a minute

OpenStudy (experimentx):

OH ... what do you mean by trigonometric form?? if it's in the from of cos \theta + i sin \theta then just subtract angle from 90 ... on both cos and sin

OpenStudy (maheshmeghwal9):

\[\cos(360^ok+58^o)+i \sin (360^ok+58^o), k \in z\]

OpenStudy (maheshmeghwal9):

why should i do that @experimentX Plz tell

OpenStudy (experimentx):

Hmm ... trigonometric functions are periodic in 360 ... adding or subtracting 360 or it's multiple does not have any effect. so ti's general solution. plus 90 - 32 = 58 and sin(90 - \theta) = cos \theta

OpenStudy (shubhamsrg):

ahh i see.. the book wants you to answer like this : substitute cos 32 = sin 58 and sin 42 = cos58 also note that since periodicity of both these func is 2pi,,you may easily write cos(58) = cos(2npi + 58) and similar with sin(58) then you'll get the ans which book says..

OpenStudy (maheshmeghwal9):

k! thanx @shubhamsrg & @experimentX :)

OpenStudy (experimentx):

yw

OpenStudy (shubhamsrg):

glad to help ^_^

OpenStudy (maheshmeghwal9):

^_^